ERUISES PPLEMENTARY EXERCISES Applications 9 In the past a chemical plant has produced an average of 1,100 Ib of chemical per day. A random sample of 260 operating days from the past year shows that x = 1,040,im1gh Man The plant manager wishes to test whether the average daily production i has dropped significantly over the past year. . Do the data provide sufficient evidence to indicate a drop in average daily production?
Please show work on how you get the answer.
Answer #1

Given : n = 260 \bar{X}=1040s^2=130000 Therefore , s = 360.5551   \mu_0=1100   a-0.05

Hypothesis : H_0:\mu=1100   Vs   H_a:\mu<1100

The test statistic under Ho is ,

Z=\frac{\bar{X}-\mu_0}{s/\sqrt{n}}\to N(0,1)

=\frac{1040-1100}{360.5551/\sqrt{260}}=-2.6833

Critical value : Z_{\alpha}=Z_{0.05}=1.64

Decision Rule : Here , Z_{stat}=-2.6833<-Z_{\alpha}=-1.64

Therefore , reject Ho at 5% level of significance

Conclusion : Hence , there is sufficient evidence to indicate a drop in average daily production.